200.0 g of Ca(C2H3O2)2 was found in the chemistry laboratory. 

      Calculate

          i. the molecular mass of 1 mole Ca(C2H3O2) 2.

         ii. the number of  moles of Ca(C2H3O2)2  that exists   

        iii. the number of Ca(C2H3O2)2 molecules present

        iv.  the volume at STP  if Ca(C2H3O2)2 is converted to gas.

        v.  the molarity of a solution made by dissolving 200.0 g of      

                Ca(C2H3O2)2 in 200 ml of water

       


1
Expert's answer
2022-05-16T17:05:03-0400

I. M[Ca(C2H3O2)2] = 40.078 + (12.0107 × 2 + 1.00794 × 3 + 15.9994 × 2) × 2 = 158.166 g mol-1

II. 1 mol Ca(C2H3O2)has a mass of 158.166 g

200 g Ca(C2H3O2)2 is (200 g Ca(C2H3O2)2) × (1 mol Ca(C2H3O2)2 / 158.166 g Ca(C2H3O2)2) = 1.2645 mol Ca(C2H3O2)2

III. 1 mole of any substance contains 6.022×1023 atoms/molecules

N[Ca(C2H3O2)2] = (1.2645 mol) × (6.022×1023 molecules / 1 mol) = 7.615×1023 molecules

IV. At STP, 1 mole of any gas occupies a volume of 22.4 L

thus 1.2645 mol of Ca(C2H3O2)2 occupies: (1.2645 mol × 22.4 L) / 1 mol = 28.325 L of Ca(C2H3O2)2

V. Molarity of Ca(C2H3O2)2 = Moles of Ca(C2H3O2)2 / Solution volume

1.2645 mol of Ca(C2H3O2)has a mass of 200.0 g (according to the (ii) calculation)

Solution volume = 200 mL = 0.2L, CM Ca(C2H3O2)2 = 1.2645 mol / 0.2 L = 6.3225 mol/L = 6.3225 M